OpenStudy (anonymous):

If f(1)=3 and f(x+1)=f(x)+2, what is f(5)?

OpenStudy (anonymous):

f(1) = 3 f(2) = f(1) + 2 = 3 + 2 = 5 f(3) = f(2) + 2 = 5 + 2 = 7 f(4) = f(3) + 2 = 7 + 2 = 9 f(5) = ...see pattern......= 11 f(6) = ..............................13

OpenStudy (anonymous):

what if you were asked to find f(12345), would you still go by the above pattern?

OpenStudy (anonymous):

I think one will need to find the equation here first and then plug in x....

OpenStudy (anonymous):

2x12345 + 1

OpenStudy (anonymous):

f(x) = 2x + 1